1.Prove that for real x, tan(x+a)/tan(x-a) when x≠∏/4, 3∏/4 cannot lie between (1-sin2a)/(1+sin2a) and (1+sin2a)/(1-sin2a)


2.Find the set of possible values of tan(x+∏/6)/tanx  where x is any real angle.

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear manoj

let

    y = tan(x+a) /tan(x-a)

       = sin(x+a)cos(x-a) / cos(x+a) sin(x-a)

       =[sin2x + sin2a]/[sin2x -sin2a]

 use componendo and divedendo

(y+1)/(y-1) = sin2x /sin2a

or

   sin2x = (y+1)sin2a /(y-1)

for every real value of x

-1≤  sin2x ≤1

 so  -1≤  (y+1)sin2a /(y-1) ≤1

 simplyfy each side of inequality

  (1-sin2a)/(1+sin2a)  ≤  y  ≤   (1-sin2a)/(1+sin2a)

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

Now you can win exciting gifts by answering the questions on Discussion Forum. So help

discuss any query on askiitians forum and become an Elite Expert League askiitian.